# Byzantine Resilient Federated Multi-Task Representation Learning

### v1

@@ Abstract
In this paper, we propose BR-MTRL, a Byzantine-resilient multi-task representation learning framework that handles faulty or malicious agents. Our approach leverages representation learning through a shared neural network model, where all clients share fixed layers, except for a client-specific final layer. This structure captures shared features among clients while enabling individual adaptation, making it a promising approach for leveraging client data and computational power in heterogeneous federated settings to learn personalized models. To learn the model, we employ an alternating gradient descent strategy: each client optimizes its local model, updates its final layer, and sends estimates of the shared representation to a central server for aggregation. To defend against Byzantine agents, we employ geometric median aggregation for robust client-server communication. Our method enables personalized learning while maintaining resilience in distributed settings. We implemented the proposed alternating gradient descent algorithm in a federated testbed built using Amazon Web Services (AWS) platform and compared its performance with various benchmark algorithms and their variations. Through extensive experiments using real-world datasets, including CIFAR-10 and FEMINIST, we demonstrated the effectiveness and robustness of our approach and its transferability to new unseen clients with limited data, even in the presence of Byzantine adversaries.
@@ I Introduction
Machine learning has traditionally relied on centralized settings, where a single model is trained on large datasets stored in one central location. This approach requires gathering data from all clients into a central data center, which raises significant data privacy concerns. Federated Learning (FL) addresses these challenges by using a distributed framework that preserves data privacy while overcoming the problem of isolated data sources [1]. FL enables multiple clients, such as mobile devices or institutions, to collaboratively train a model without sharing their raw data. A central server is responsible for coordinating the clients by aggregating the models received from each client and broadcasting the updated model back to them in each round of communication. McMahan et al. introduced the Federated Averaging (FedAvg) algorithm, which facilitates the training of a single shared global model designed to achieve optimal performance on average across all participating clients [1]. The approach offers several advantages, such as simplicity in implementation and a streamlined aggregation process, making it scalable for large client populations. Additionally, the global model promotes uniformity, ensuring consistent performance for clients with similar data distributions. However, FL approaches encounter three primary bottlenecks: (i) Data heterogeneity: In real-world scenarios, client tasks often exhibit significantly different underlying data distributions. In such environments, a global model optimized for average loss may underperform across diverse clients, as it fails to account for these variations in data [2]. (ii) Generalization and adaptability: Federated models may struggle to generalize to unseen clients that were not part of the training process, limiting their effectiveness in dynamic, real-world applications [3]. (iii) Byzantine attacks: FL is particularly vulnerable to Byzantine attacks, as it relies on distributed client updates; a faulty or malicious client can potentially corrupt the global model, compromising performance and reliability [4].
@@
Despite these advancements, a unified approach for Byzantine-resilient personalized FL with strong generalization remains unavailable. In this paper, we introduce the Byzantine-resilient Multi-Task Representation Learning (BR-MTRL) framework to jointly address three critical challenges in FL: data heterogeneity, generalization, and Byzantine resilience. By incorporating geometric median aggregation methods into personalized FL, we enhance the global model’s resilience against Byzantine attacks. Our framework ensures that the models learned across distributed agents remain robust, reliable, and secure, even in the presence of adversarial influences, improving both performance and safety.
@@ I-A Contributions
  • We propose a Byzantine-resilient MTRL algorithm based on personalized federated learning. In our approach, all clients share a common representation while maintaining their local personalized heads. We introduce an alternating gradient descent approach that alternates between solving for the shared model and fine-tuning each client’s personalized head. To ensure robustness, our method employs a Geometric Median (GM)-based mechanism for aggregating client updates at the server, allowing the clients to learn a Byzantine-resilient shared representation while preserving task-specific personalizations.
@@ I-B Related Work
Personalized FL has received a lot of research interest in recent years. Liang et al. [16] proposed the Local Global Federated Averaging (LG-FedAvg) method to address data heterogeneity, which focuses on learning low-dimensional features from clients’ data. Their approach emphasizes learning multiple local representations and a single global head instead of using a single global representation and multiple local heads. On the other hand, Arivazhagan et al. [17] proposed FedPer algorithm that explicitly divides the model architecture into base layers, trained across all devices via federated averaging and personalization layers, which are tailored locally on each device using its specific data. This structure not only promotes personalization by adapting the model to local data characteristics but also maintains a cohesive model performance across diverse environments by leveraging a shared base learning across all clients. In addition, Collins et al. [3] proposed the FedRep algorithm focusing on learning a global model shared across clients while allowing each client to maintain a personalized head model, similar to [17]. The algorithms in [3, 17] allow clients to collaborate on learning shared representations while maintaining personalized models tailored to their specific data distribution. [3, 17] are closely related to our work; however, they focus on personalized FL without considering the challenges posed by Byzantine attacks.
@@
To counter Byzantine attacks, many personalized FL systems incorporate robust aggregation techniques at the server, such as the geometric median and variations of the geometric median. For example, Li et al. introduced AutoGM, an auto-weighted geometric median aggregation rule designed to enhance robustness in the presence of Byzantine failures, specifically model and data poisoning attacks [18]. AutoGM improves the classic geometric median by automatically excluding extreme outliers and re-weighting the remaining parameter updates based on a skewness threshold, offering a more adaptive and flexible solution. Similarly, Wang et al. proposed WGM-dSAGA, a federated learning strategy that employs a weighted geometric median for robust aggregation against Byzantine attacks [19]. The strategy incorporates COPOD, a parameter-free outlier detection technique, to identify and assign weights to client updates based on their potential maliciousness. This method effectively mitigates the influence of Byzantine clients on the global model through the weighted geometric median. Meanwhile, He et al. tackled Byzantine attacks by introducing a novel projected stochastic block gradient descent method [20]. Unlike the GM-based approaches, their strategy employs a server aggregation technique that combines local models using a Huber function-based descent step to manage the impact of malicious updates from Byzantine clients. While these studies have made significant strides in mitigating Byzantine attacks, they primarily focus on enhancing aggregation techniques on the server side. In contrast, our work introduces a Byzantine-resilient MTRL framework that learns and leverages a shared representation to enhance the quality of each client’s model, while ensuring both transferability and resilience.
@@ II-A System Model
Robust server. To minimize the impact of Byzantine clients, we apply the Geometric Median (GM) to the weights received from clients, which is then transmitted to the server. GM offers a more resilient aggregation method by effectively reducing the influence of outlier updates. By using the median instead of the mean, this approach reduces outlier distortion, ensuring the global model reflects honest client updates.
@@ II-B Design Goals
Byzantine robustness. The proposed scheme learns an unbiased personalized models, which minimizes the impact of malicious model updates.
@@ III Proposed Byzantine Resilient MTRL Algorithm
Client Update. In each round $t\in[T]$, we select a fraction $r\in(0,1]$ of clients to participate in the client update. During this update, client $i$ performs ${\tau_{h}}$ local gradient steps to optimize its head model based on the current global representation ${\phi^{t-1}}$, which is shared by the server. Specifically, for each step $s=1,\dots,{\tau_{h}}$, client $i$ updates its head model as follows:
@@
10:         for $s=1$ to $\tau_{\phi}$ $\triangleright$ Client $i$ makes $\tau_{\phi}$ shared model updates
@@
21:     Server computes the new representation as $\phi^{t}\leftarrow\mathrm{GM}(\phi_{1}^{t,\tau_{\phi}},\cdots,\phi_{n}^{t,\tau_{\phi}})$
@@ IV-A Experimental Setup
Benchmark Algorithms. We evaluated our algorithm with benchmark approaches, including FedAvg [1], FedPer [17], and FedRep [3]. FedPer and FedRep are similar approaches that learn a shared representation with personalized heads, but FedPer updates both simultaneously, maintaining an equal number of updates per local round. In contrast, FedAvg learns a single global model without personalization. As a best-case benchmark, we considered these methods without malicious clients. Additionally, we analyzed two variants: (i) introducing malicious clients (Byzantine FedPer/FedRep/FedAvg) and (ii) incorporating geometric median-based aggregation (FedAvg+GM, FedPer+GM). For meta-test, we consider an additional baseline, Naive, which trains new clients independently without leveraging the MTRL framework.
@@ IV-C Experiment Results
Figure 2:   FedRep,   Byzantine FedRep,   BR-MTRL (proposed algorithm),   Naive. Average test accuracy across all clients: Figure 2(a) shows the results for CIFAR-10 with 100 clients, including 20 malicious clients. Figure 2(b) presents the results for FEMNIST with 150 clients, including 20 malicious clients. We compared our proposed BR-MTRL with the FedRep algorithm under benign conditions (i.e., no malicious client) and FedRep subjected to Byzantine attacks (Byzantine FedRep). Meta test for new clients: Figures 2(c) and 2(d) present the average test accuracy across 10 new clients (CIFAR-10 and FEMNIST) after fine-tuning the local heads for 10 epochs using the learned shared model. We first learned the shared model using a set of source clients and then used the shared model to optimize the local head of new clients.
@@
BR-MTRL (proposed) 72.94 58.55 75.92 55.28
@@
Test accuracy of BR-MTRL. As shown in Figures 2(a) and 2(b) and Table II, the FedRep and FedPer algorithms consistently achieved higher accuracy than FedAvg across various experiments, irrespective of the number of clients or classes per client, demonstrating the effectiveness of personalized FL. However, the performance of these three baselines (best-case without considering malicious client), significantly declined under Byzantine attacks. This decline validates that (personalized) FL frameworks susceptible to Byzantine attacks will lead to unreliable models with poor performance. Increasing the number of classes per client from two to five resulted in a substantial accuracy drop of $21.61\%$. This decline occurs because Byzantine clients can negatively impact the global model’s performance across a greater number of classes.
@@
Furthermore, reducing the proportion of Byzantine clients relative to the total number of clients improves performance, as seen in the comparison between experiments with $(100,2,20)$ and $(1000,2,100)$, i.e., $20\%,10\%$ malicious clients, respectively, where the latter exhibits greater resilience. Our experimental results further demonstrate that the proposed approach achieves a higher accuracy than the GM-adapted version of FedPer. This is because FedPer updates the server model at the same frequency as client updates, making it more susceptible to Byzantine clients, which can adversely affect learning accuracy. In contrast, our approach updates the shared model only once after updating the client heads, leading to improved overall performance. All experimental results consistently showcase the improved performance of our proposed approach.
@@
Results for Meta-test. In this experiment, we tested the transferability of the proposed approach. We first learn the shared model $\phi$ using the source clients and then utilize this model to train the local heads of new, previously unseen clients. For each new client, we fine-tune the client head for $10$ epochs, using the transferred shared model. For CIFAR-10, we randomly selected 10 clients as new clients and trained the shared model using the remaining 90 clients, including 20 Byzantine clients. Similarly, for FEMNIST, we selected 10 new clients and trained the shared model with the remaining 140 clients, of which 10 were Byzantine. We evaluated the average accuracy for four approaches. To highlight the benefits of MTRL for transfer learning, we considered a Naive setting, where each client trains their model independently, resulting in poor accuracy. Figures 2(c) and 2(d) illustrate the results: FedRep, without Byzantine attacks, demonstrates effective transferability of the shared model to new clients. In contrast, Byzantine MTRL, with malicious clients, exhibits significantly lower performance, highlighting the vulnerability of FL systems. Finally, BR-MTRL, our proposed approach, improves learning accuracy by employing GM aggregation, validating its effectiveness.
@@ V Conclusion
In this paper, we proposed a Byzantine-resilient federated MTRL framework to learn robust shared representations across related yet different tasks. Our approach employs an alternating gradient descent strategy, where the shared model is updated while keeping the local head fixed, and vice versa. To ensure Byzantine resilience, we incorporated geometric median-based aggregation at the server, mitigating the impact of malicious clients sending corrupted updates. We implemented our federated learning approach on AWS, simulating a real-world federated setting. Extensive experiments on CIFAR-10 and FEMNIST validate the effectiveness of our approach and its transferability to new clients with limited data.

### v2

@@ Abstract
In this paper, we propose BR-MTRL, a Byzantine-resilient multi-task representation learning framework that handles faulty or malicious agents. Our approach leverages representation learning through a shared neural network model, where all clients share fixed layers, except for a client-specific final layer. This structure captures shared features among clients while enabling individual adaptation, making it a promising approach for leveraging client data and computational power in heterogeneous federated settings to learn personalized models. To learn the model, we employ an alternating gradient descent strategy: each client optimizes its local model, updates its final layer, and sends estimates of the shared representation to a central server for aggregation. To defend against Byzantine agents, we employ two robust aggregation methods for client-server communication, Geometric Median and Krum. Our method enables personalized learning while maintaining resilience in distributed settings. We implemented the proposed algorithm in a federated testbed built using Amazon Web Services (AWS) platform and compared its performance with various benchmark algorithms and their variations. Through experiments using real-world datasets, including CIFAR-10 and FEMNIST, we demonstrated the effectiveness and robustness of our approach and its transferability to new unseen clients with limited data, even in the presence of Byzantine adversaries.
@@ I Introduction
Machine learning has traditionally relied on centralized settings, where a single model is trained on large datasets stored in one central location. This approach requires gathering data from all clients into a central data center, which raises significant data privacy concerns. Federated Learning (FL) addresses these challenges by using a distributed framework that preserves data privacy while overcoming the problem of isolated data sources [1]. FL enables multiple clients, such as mobile devices or institutions, to collaboratively train a global model without sharing their raw data. A central server is responsible for coordinating the clients by aggregating the models received from each client and broadcasting the updated model back to them in each round of communication. McMahan et al. introduced the Federated Averaging (FedAvg) algorithm, which facilitates the training of a single shared global model designed to achieve optimal performance on average across all participating clients [1]. The approach offers several advantages, such as simplicity in implementation and a streamlined aggregation process, making it scalable for large client populations. Additionally, the global model promotes uniformity, ensuring consistent performance for clients with similar data distributions. However, FL approaches encounter three primary bottlenecks: (i) Data heterogeneity: In real-world scenarios, client tasks often exhibit significantly different underlying data distributions. In such environments, a global model optimized for average loss may underperform across diverse clients, as it fails to account for these variations in data [2]. (ii) Generalization and adaptability: Federated models may struggle to generalize to unseen clients that were not part of the training process, limiting their effectiveness in dynamic, real-world applications [3]. (iii) Byzantine attacks: FL is particularly vulnerable to Byzantine attacks, as it relies on distributed client updates; a faulty or malicious client can potentially corrupt the global model, compromising performance and reliability [4].
@@
Despite these advancements, a unified approach for Byzantine-resilient personalized FL with strong generalization remains unavailable. In this paper, we introduce the Byzantine-resilient Multi-Task Representation Learning (BR-MTRL) framework to jointly address three critical challenges in FL: data heterogeneity, generalization, and Byzantine resilience. By incorporating geometric median and Krum aggregation methods into personalized FL, we enhance the global model’s resilience against Byzantine attacks. Our framework ensures that the models learned across distributed agents remain robust, reliable, and secure, even in the presence of adversarial influences, improving both performance and safety.
@@ I-A Contributions
  • We propose a Byzantine-resilient MTRL algorithm based on personalized federated learning. In our approach, all clients share a common representation while maintaining their local personalized heads. We introduce an alternating gradient descent approach that alternates between solving for the shared model and fine-tuning each client’s personalized head. To ensure robustness, our method employs Geometric Median (GM) and Krum-based mechanisms for aggregating client updates at the server, allowing the clients to learn a Byzantine-resilient shared representation while preserving task-specific personalizations.
@@ I-B Related Work
Personalized Federated Learning (PFL) has received a lot of research interest in recent years. Liang et al. [16] proposed the Local Global Federated Averaging (LG-FedAvg) method to address data heterogeneity, which focuses on learning low-dimensional features from clients’ data. Their approach emphasizes learning multiple local representations and a single global head. On the other hand, Arivazhagan et al. [17] proposed FedPer algorithm that explicitly divides the model architecture into base layers, trained across all devices via federated averaging and personalization layers, which are tailored locally on each device using its specific data. This structure not only promotes personalization by adapting the model to local data characteristics but also maintains a cohesive model performance across diverse environments by leveraging a shared base learning across all clients. In addition, Collins et al. [3] proposed the FedRep algorithm focusing on learning a global model shared across clients while allowing each client to maintain a personalized head model, similar to [17]. The algorithms in [3, 17] allow clients to collaborate on learning shared representations while maintaining personalized models tailored to their specific data distribution. [3, 17] are closely related to our work; however, they focus on PFL without considering the challenges posed by Byzantine attacks.
@@
To counter Byzantine attacks, many PFL systems incorporate robust aggregation techniques at the server, such as the geometric median and its variations. For example, Li et al. introduced AutoGM, an auto-weighted geometric median aggregation rule designed to enhance robustness in the presence of Byzantine failures, specifically model and data poisoning attacks [18]. AutoGM improves the classic geometric median by automatically excluding extreme outliers and re-weighting the remaining parameter updates based on a skewness threshold, offering a more adaptive and flexible solution. Similarly, Wang et al. proposed WGM-dSAGA, a federated learning strategy that employs a weighted geometric median for robust aggregation against Byzantine attacks [19]. The strategy incorporates COPOD, a parameter-free outlier detection technique, to identify and assign weights to client updates based on their potential maliciousness. This method effectively mitigates the influence of Byzantine clients on the global model through the weighted geometric median. Meanwhile, He et al. tackled Byzantine attacks by introducing a novel projected stochastic block gradient descent method [20]. Unlike the GM-based approaches, their strategy employs a server aggregation technique that combines local models using a Huber function-based descent step to manage the impact of malicious updates from Byzantine clients. While these studies have made significant strides in mitigating Byzantine attacks, they primarily focus on enhancing aggregation techniques on the server side. In contrast, our work introduces a Byzantine-resilient MTRL framework that learns and leverages a shared representation to enhance the quality of each client’s model, while ensuring both transferability and resilience.
@@ II-A System Model
Robust server. To minimize the impact of Byzantine clients, we apply the Geometric Median (GM) and Krum to the weights received from clients, which is then transmitted to the server. These techniques offer a more resilient aggregation by effectively reducing the influence of outlier updates. By using the median instead of the mean, GM reduces outlier distortion, ensuring the global model reflects honest client updates. Krum enhances resilience by selecting the client update that is most similar to its nearest neighbors, thereby mitigating the impact of malicious updates on the global model.
@@ II-B Design Goals
Byzantine robustness. The proposed scheme learns an unbiased, personalized model, which minimizes the impact of malicious model updates.
@@ III Proposed Byzantine Resilient MTRL Algorithm
Client Update. In each round $t\in[T]$, we select a fraction $\alpha\in(0,1]$ of clients to participate in the client update. During this update, client $i$ performs ${\tau_{h}}$ local gradient steps to optimize its head model based on the current global representation ${\phi^{t-1}}$, which is shared by the server. Specifically, for each step $s=1,\dots,{\tau_{h}}$, client $i$ updates its head model as follows:
@@
Further, we perform a Krum approach [14] detailed below.
${\mathrm{Krum}}(\phi_{1}^{t,\tau_{\phi}},\cdots,\phi_{n}^{t,\tau_{\phi}})=\phi_{k^{*}}^{t,\tau_{\phi}},\mbox{~{}where}$
$k^{*}=\arg\min_{k\in[n]}\sum_{k\rightarrow k^{\prime}}\|\phi_{k}^{t,\tau_{\phi}}-\phi_{k^{\prime}}^{t,\tau_{\phi}}\|_{F}.$
For $k\neq k^{\prime}$, $k\rightarrow k^{\prime}$ is the set of $n-f-2$ nearest neighbors of $\phi_{k}^{t,\tau_{\phi}}$, where $f$ is the number of Byzantine clients. Thus $\phi_{k^{*}}^{t,\tau_{\phi}}$ is the model with minimal summed distance to its neighbors.
Remark III.1.
Krum requires to know the number of Byzantine clients in advance, and its computational complexity scales quadratically with the number of clients, resulting in significantly higher computational requirements as compared to GM.
@@
10:         for $s=1$ to $\tau_{\phi}$ $\triangleright$ Client $i$ makes $\tau_{\phi}$ updates to $\phi$
@@
[…]
@@ IV-A Experimental Setup
Benchmark Algorithms. We evaluated our algorithm with benchmark approaches, including FedAvg [1], FedPer [17], and FedRep [3]. FedPer and FedRep are similar approaches that learn a shared representation with personalized heads, but FedPer updates both simultaneously, maintaining an equal number of updates per local round. In contrast, FedAvg learns a single global model without personalization. As a best-case benchmark, we considered these methods without malicious clients. Additionally, we analyzed two variants: (i) introducing malicious clients (Byzantine FedPer/FedRep/FedAvg), (ii) incorporating geometric median-based aggregation (FedAvg+GM, FedPer+GM), and (iii) incorporating Krum-based aggregation (FedAvg+Krum, FedPer+Krum). For meta-test, we consider an additional baseline, Naive, which trains new clients independently without leveraging the MTRL framework.
@@ IV-C Experiment Results
Figure 2:   FedRep,   Byzantine FedRep,   BR-MTRL+GM (proposed),   BR-MTRL+Krum (proposed)   Naive. Average test accuracy across all clients: Figure 2(a) shows the results for CIFAR-10 with 100 clients with 20 malicious clients. Figure 2(b) presents the results for FEMNIST with 150 clients with 20 malicious clients. We compared our proposed BR-MTRL with the FedRep algorithm under benign conditions (i.e., no malicious client) and FedRep subjected to Byzantine attacks (Byzantine FedRep). Meta test for new clients: Figures 2(c) and 2(d) present the average test accuracy across 10 new clients (CIFAR-10 and FEMNIST) after fine-tuning the local heads for 10 epochs using the shared model. We first learned the shared model using source clients and then used the shared model to optimize the local head of new clients.
@@
FedPer+Krum 71.90 61.21 68.17 53.12
BR-MTRL+GM 72.94 58.55 75.92 55.28
BR-MTRL + Krum 80.25 61.80 72.76 54.07
@@
Test accuracy. As shown in Figures 2(a) and 2(b) and Table II, the FedRep and FedPer algorithms consistently achieved higher accuracy than FedAvg across various experiments, irrespective of the number of clients or classes per client, demonstrating the effectiveness of personalized FL. However, the performance of these three baselines (without any malicious clients) significantly declined under Byzantine attacks. This decline validates that (personalized) FL frameworks susceptible to Byzantine attacks will lead to unreliable models with poor performance. Increasing the number of classes per client from two to five resulted in a substantial accuracy drop of $21.61\%$. This is because Byzantine clients can negatively impact the global model’s performance across a greater number of classes. Furthermore, reducing the proportion of Byzantine clients relative to the total number of clients improves performance, as seen in the comparison between experiments with $(100,2,20)$ and $(1000,2,100)$, i.e., $20\%,10\%$ malicious clients, respectively, where the latter exhibits greater resilience.
@@
Our experimental results further demonstrate that the proposed approaches (BR-MTRL+GM and BR-MTRL+Krum) achieves a higher accuracy than both the GM and Krum-adapted versions of FedPer. This is because FedPer updates the server model at the same frequency as client updates, making it more susceptible to Byzantine clients, which can adversely affect learning accuracy. In contrast, our approach updates the shared model only once after updating the client heads, leading to improved overall performance. All experimental results consistently showcase the improved performance of proposed approach. Krum outperforms GM in experiments with $100$ clients; however, has a lower accuracy when the number of clients increases. Krum requires to know the number of malicious clients and computational complexity is higher.
@@
Results for Meta-test. We tested the transferability of the proposed approach. We first learn the shared model $\phi$ using the source clients and then utilize this model to train the local heads of new, previously unseen clients. For each new client, we fine-tune the client head for $10$ epochs, using the transferred shared model. For CIFAR-10, we randomly selected 10 clients as new clients and trained the shared model using the remaining 90 clients, including 20 Byzantine clients. Similarly, for FEMNIST, we selected 10 new clients and trained the shared model with the remaining 140 clients, of which 10 were Byzantine. We evaluated the average accuracy for four approaches. To highlight the benefits of MTRL for transfer learning, we considered a Naive setting, where each client trains their model independently, resulting in poor accuracy. Figures 2(c) and 2(d) illustrate the results: FedRep, without Byzantine attacks, demonstrates effective transferability of the shared model to new clients. In contrast, Byzantine MTRL, with malicious clients, exhibits significantly lower performance, highlighting the vulnerability of FL systems. Finally, BR-MTRL, our proposed approach, improves learning accuracy by employing GM and Krum aggregation, validating its effectiveness.
@@ V Conclusion
In this paper, we proposed a Byzantine-resilient federated MTRL framework to learn robust shared representations across related yet different tasks. Our approach employs an alternating gradient descent strategy, where the shared model is updated while keeping the local head fixed, and vice versa. To ensure Byzantine resilience, we incorporated geometric median and Krum-based aggregation at the server, mitigating the impact of malicious clients sending corrupted updates. We implemented our federated learning approach on AWS, simulating a real-world federated setting. Extensive experiments on CIFAR-10 and FEMNIST validate the effectiveness of our approach and its transferability to new clients with limited data.